OpenStudy (anonymous):

if F(x)=-3x+7 and g(x)=-7x+3 what is the value of f(-3)-g(3)

OpenStudy (anonymous):

(-3(-3)+7)-(-7(3)+3)

OpenStudy (anonymous):

Hi welcome to Openstudy! I would suggest that it would be easier for you to substitute all the values of x in the functions F(x) and g(x). for example: F(x)=-3x+7 g(x)=-7x+3 F(-3)= -3(-3)+7 g(3)=-7(3) +3 then after you solve for the values of x from both functions, subtract the value of \(g\) from the value of \(f\)

OpenStudy (anonymous):

thats what i thought but the answer choices are 40,34,8, or -8

OpenStudy (alesia/bailey):

i got like -26 if i did it right

OpenStudy (anonymous):

how did you get -26 can you show

OpenStudy (anonymous):

the answer i got is one of those choices

OpenStudy (anonymous):

ill just go with 40 then thanks

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it's not 40 either

OpenStudy (anonymous):

when i solved it my answer wasn't a choice

OpenStudy (anonymous):

show your work please ill show you where you messed up

OpenStudy (anonymous):

i solved the equation that you sent before and i got 40, then i solved what the next person said to do and my answer wasn't a chocie

OpenStudy (anonymous):

i didnt get 40 when i solved

OpenStudy (anonymous):

what did you get, i just solved it again and still got 40

OpenStudy (anonymous):

you're right im mistake was on my end

OpenStudy (anonymous):

im sorry mistake was on my end*